The hypotenuse of a right triangle is twice as long as one leg and 20 cm longer than the other leg. The perimeter is 240 cm. What is the area of the triangle?
Answer by ewatrrr(24785)

x the length of the hypotenuse
x + x/2 + (x-20) = 240 |Multiplying thru by 2 so as all denominators = 1
2x + x + 2(x-20) = 480
5x = 520
x = 104, length of the hypotenuse. Legs: 52 and 84
And... 104 + 52 + 84 = 240
A = (1/2)52•84 = 2184 cm^2